OpenStudy (anonymous):

Find the midpoint between the two points: (2,2), (6,4) i got (4,3) @sangya21

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

midpoint of x = (x1+x)/2 midpoint of y = (y1+y)/2

OpenStudy (anonymous):

ok